40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 27 20:41:21 2017, rana, gautam, Update, PSL, PMC DAQ assay for feed-forward integration TEK00000.PNG20170327_194931.jpg20170327_204554.jpg
    Reply  Wed Mar 29 11:41:19 2017, gautam, Update, PSL, PMC DAQ assay for feed-forward integration C1ALS_PSLnamespace.pngPMCcheckout.pdfD980352-A-40m_151119.pdf
       Reply  Thu Mar 30 00:16:09 2017, gautam, Update, PSL, PSL NPRO PZT calibration 
          Reply  Mon Apr 3 17:25:13 2017, gautam, Update, PSL, PSL NPRO PZT calibration PDHerr.pdfNPROcalib.pdf
             Reply  Tue Apr 4 17:27:58 2017, rana, Update, PSL, PSL NPRO PZT calibration 
       Reply  Thu Mar 30 18:11:01 2017, rana, Update, PSL, PMC DAQ assay for feed-forward integration 
          Reply  Mon Apr 3 23:07:09 2017, gautam, Update, PSL, PMC DAQ assay for feed-forward integration 
             Reply  Tue Apr 11 00:38:37 2017, gautam, Update, PSL, PMC demod moved off servo board PMC_Ctrl_spec.pdf
                Reply  Wed Apr 12 00:36:53 2017, gautam, Update, PSL, PMC demod moved off servo board PMC_digitalSpec.pdfPMC_DAQ_spectra.pdf
                   Reply  Tue Apr 18 01:01:03 2017, gautam, Update, PSL, PMC OLTF measured, DAQ channels calibrated PMCspectra_calibrated.pdfPMC_MEDMs.png
                      Reply  Tue Apr 18 16:10:00 2017, gautam, Update, PSL, PMC OLTF measured, DAQ channels calibrated PMC_OLTF_170418.pdf
                         Reply  Tue Apr 18 23:37:15 2017, rana, Update, PSL, PMC OLTF measured, DAQ channels calibrated 
                         Reply  Wed Apr 19 15:13:30 2017, gautam, Update, PSL, PMC/MCL multicoherence PMC_MCL_multicoherence.pdf
Message ID: 12926     Entry time: Mon Apr 3 23:07:09 2017     In reply to: 12920     Reply to this: 12939
Author: gautam 
Type: Update 
Category: PSL 
Subject: PMC DAQ assay for feed-forward integration 

I made some changes to the DAQ path on the PMC servo board, as per the plan posted earlier in this thread. Summary of changes:

  1. AC coupling PMC control signal path using 2 x 47uF metal film capacitors (in parallel)
  2. Grounding pin 5 of U15
  3. Adding gain to U14 (gain of ~500) and U15 (gain of ~50)

Details + photos + calibration of DAQ channels to follow. The PMC and IMC both seem to remain stably locked after this work.

ELOG V3.1.3-